Sol Summary Report – March 6th

Crew 275 Sol Summary Report 06Mar2023

Sol: 22

Summary Title: “The last week’s a smooth ride”, they said!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Another 4-hour EVA left three crewmembers very tired today. They went to Candor Chasma and successfully performed the photogrammetry EVA, based on a 3D render of the region. Only one crewmember could take part as a subject, since the other subject was injured. Before reaching Candor Chasma, they stopped to install the LOAC, and then to change the batteries and retrieve the data of the other atmospheric instruments. They noticed that MegaAres was off, as it ran out of battery. We have noticed that about 3 hours of recording are missing. This is due to the fact that we did not have enough time and enough chargers to completely charge the batteries of all the atmospheric instruments.

With the GreenHab officer in EVA, the Crew Astronomer watered the crops. We then spent some additional time in the GreenHab to perform a few cognitive tests, while the temperature there was very close to 100°F/40°C.

The injured crewmember is feeling slightly better, and we managed to get an appointment with a surgeon for her. We will drive her to Richfield tomorrow. However, the fact that a few crewmembers will not be at the MDRS required us to modify the schedule and to perform additional activities today.

Look Ahead Plan:

Today, we managed to get an appointment with a surgeon in Richfield. The injured crewmembers will have to be at the Sevier Valley hospital tomorrow (Sol 23) at 10:40 for a second X-ray and then to see the surgeon. We requested Mission Support to use the CrewCar to do so. The plan is to leave between 7:00 and 7:30 with the Crew Journalist, the Crew Scientist, and the Health & Safety Office, so that two drivers can accompany the injured crewmember. In the absence of the Health & Safety Officer, the Commander will take this role. The tasks of the Crew Scientist and of the Crew Journalist will be shared between the four crewmembers remaining at the MDRS.

A media visit is planned for tomorrow (Sol 23) with a crew from France Televisions. They will arrive at 9:00 and leave at 17:00. In the morning, we plan to have them following us during our EVA, first to change the battery and retrieve the data from the atmospheric instruments, and then to shoot nice videos at North Ridge. In the afternoon, we will show them the different buildings and the rest of our experiments.

We would also like to perform an exploration EVA on Thursday (Sol 25) and another EVA dedicated to the disassembling of the atmospheric instruments on Friday (Sol 26).

Anomalies in work: None.

Weather: Windy, low 1°C/34°F, high 12°C/54°F.

Crew Physical Status: One crewmember injured (broken collarbone / clavicula). Appointment with surgeon is planned tomorrow (Sol 23) at 11:10 in Richfield. Surgery should soon follow.

EVA: EVA # 18 dedicated to the photogrammetry experiment based on a 3D render was successfully performed. Some anomalies were observed regarding the batteries of the atmospheric instruments.

Reports to be filed: Sol Summary, EVA Report, EVA Request, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site